OpenStudy (anonymous):

For the triangle with vertices P(-2, 0), Q(4, -6), and R(5, -3), find an equation for the median from vertex Q.

OpenStudy (anonymous):

Median bisects the opposite side, so first you need to find that side's midpoint. Then, find the line that passes through those two points. Do you need help with that part?

OpenStudy (anonymous):

Yes please. I have the midpoint of PR which is (1.5, -1.5). Do I have to use y=mx+b?

OpenStudy (anonymous):

Yes. Do you know how to find slope?

OpenStudy (anonymous):

m=rise/run?

OpenStudy (anonymous):

Yeah. What's the slope?

OpenStudy (anonymous):

Is it 7.5/2.5? I kind of suck at slope.

OpenStudy (anonymous):

Rise is difference in y and run is difference in x. Q(4.0, -6.0) M(1.5, -1.5) Difference in y: -6 - -1.5 = -4.5 Difference in x: 4 - 1.5 = 2.5 Slope m = -4.5/2.5 = -9/5

OpenStudy (anonymous):

Shoot! I just realised I typed the question wrong. Q is (4, 6) not (4, -6). Sorry

OpenStudy (anonymous):

No problem, it's the same process.

OpenStudy (anonymous):

Slope should be 3 (typed it in a slope calculator).

OpenStudy (anonymous):

Okay , I see how you got 3 because 7.5/2.5 simplified is 3.

OpenStudy (anonymous):

Yeah. Now all you need to do is find b in y = mx + b. You have m, and you have x and y (pick a point that you know is on the line, either the midpoint or Q). Plug everything in and solve for b. Then you have m and you have b, so set up your answer as y = mx + b and you're done.

OpenStudy (anonymous):

So is my final answer y=3x+(-6)?

OpenStudy (anonymous):

I think so! I used http://www.basic-mathematics.com/slope-calculator.html and that's what it gave me.

OpenStudy (anonymous):

Okay. Thank you so much!